### v1

# Exploiting the Layered Intrinsic Dimensionality of Deep Models for Practical Adversarial Training

Abstract
Despite being a heavily researched topic, Adversarial Training (AT) is rarely, if ever, deployed in practical AI systems for two primary reasons: (i) the gained robustness is frequently accompanied by a drop in generalization and (ii) generating adversarial examples (AEs) is computationally prohibitively expensive. To address these limitations, we propose SMAAT  a new AT algorithm that leverages the manifold conjecture [49], stating that off-manifold AEs lead to better robustness while on-manifold AEs result in better generalization. Specifically, SMAAT aims at generating a higher proportion of off-manifold AEs by perturbing the intermediate deepnet layer with the lowest intrinsic dimension. This systematically results in better scalability compared to classical AT as it reduces the PGD chains length required for generating the AEs. Additionally, our study provides, to the best of our knowledge, the first explanation for the difference in the generalization and robustness trends between vision and language models, i.e., AT results in a drop in generalization in vision models whereas, in encoder-based language models, generalization either improves or remains unchanged. We show that vision transformers and decoder-based models tend to have low intrinsic dimensionality in the earlier layers of the network (more off-manifold AEs), while encoder-based models have low intrinsic dimensionality in the later layers. We demonstrate the efficacy of SMAAT  on several tasks, including robustifying (i) sentiment classifiers, (ii) safety filters in decoder-based models, and (iii) retrievers in RAG setups. SMAAT requires only 25-33% of the GPU time compared to standard AT, while significantly improving robustness across all applications and maintaining comparable generalization.
1 Introduction
Adversarial training (AT) has been the most effective approach for training deep neural network models to be robust to small input perturbations [6, 28]. It is formulated as a min-max optimization problem where the outer minimization is over the parameter space and the inner-maximization searches for the worst-case input space perturbation. For deepnets, the inner maximization is classically solved approximately using several iterations of the projected gradient descent method (PGD, Madry et al. [38]). However, AT suffers from at least two major limitations. First, the impact of improving robustness on generalization is not well understood [55]. For instance, in vision models, AT leads to a drop in generalization (e.g., 8% drop on CIFAR10 [48]), whereas in encoder-based language models (enc-LLMs) [^1] We will use ”enc-LLMs” to refer to encoder-based language models such as BERT and RoBERTa, and ”dec-LLMs” for decoder-based language models such as LLaMA, it either improves generalization or remains unchanged (e.g., 1% improvement on AGNEWS [80]). Moreover, robustness improvement is consistently more significant in vision (e.g., 40% on CIFAR10) than in enc-LLMs (e.g., 10% on AGNEWS). While there has been a lot of effort to understand the drop in generalization in vision models [38, 65, 3, 73, 13], there has not been any work understanding the inherent difference in the robustness-generalization relationship in vision, dec-LLMs and enc-LLMs. Second, AT is associated with a high computational overhead as it requires several forward and backward passes through the network. This significantly limits AT scalability [46] to increasingly wider and deeper models [15, 8]. To address this, recent research efforts focused on minimizing the number of PGD steps, often through reusing or accumulating gradients during parameter updates [48, 72, 80]. Yet, these methods still require full-length network passes and, despite improvement, continue to be computationally demanding.
(a) Classification
(b) RAG
(c) Safety Filter
Figure 1: Comparison of SMAAT robustness, generalization, and run time (marker size) against baselines for robustifying (a) topic classifiers, (b) retriever models in the Retrieval Augmented Generation (RAG) setup and (c) safety filters for decoder-based LLMs. SMAAT significantly enhances model robustness compared to seven different baselines, while maintaining nearly the same clean accuracy. Additionally, it greatly improves the scalability of adversarial training.
In this work, we first identify a root cause of the difference in the robustness-generalization trends in foundational models. We show that the intrinsic dimension (ID) of the data manifold in the first layer is much higher in enc-LLMs compared to vision models/dec-LLM (see Figure 2) which results in AEs being more on-manifold (ONM-AEs) in enc-LLMs and more off-manifold (OFM-AEs) in vision models and dec-LLMs. In accordance with the manifold conjecture [18, 50, 21], we find that OFM-AEs lead to better robustness while more ONM-AEs lead to better generalization. To the best of our knowledge, this is the first explanation for the difference in the robustness magnitudes across vision models and enc-LLMs. Our findings are also consistent with YOPO [72] and TMD [40]. Specifically, YOPO highlights the critical role of the first layer in vision models for AT, while TMD uses the last layer to detect AEs in enc-LLMs.
To address the second limitation, i.e., scalability, we leverage our aforementioned insights on the impact of the manifold ID on robustness and generalization and hypothesize that perturbing the intermediate layer $l$ with the highest off-manifold AE ratio (equivalently lowest ID) should achieve high robustness at low computational cost. Intuitively, a lower ID corresponds to a higher proportion of OFM-AEs generation and a greater improvement in robustness by the manifold conjecture. Additionally, layers closer to the output result in shorter PGD chains, leading to efficient AT. To this end, we propose SMAAT, a Scalable Manifold Aware AT approach that applies AT at the layer with the highest proportion of OFM-AEs. We found that this critical layer is consistently the last layer across enc-LLMs on several applications. Hence, for enc-LLMs, SMAAT leads to a significant speed-up of AT by avoiding a full backward pass as it calculates the gradients only until the last layer rather than the entire network. Also, empirically, we show that we achieve a higher robustness. Yet, this ID trend is different for vision or dec-LLMs, where we find that the first layer is always the one with the lowest ID. So, SMAAT in this case would be equivalent to standard adversarial training. Nevertheless, we argue that there is significant value in robustifying enc-LLMs as they are increasingly being used to augment the capabilities (e.g., retrievers in RAG setups [23]) or improve the safety (e.g., safety filters [27]) in dec-LLMs.
Figure 2: Overview. First, we study the differences in robustness and generalization trends in AT in encoder-based (enc-), decoder-based (dec-) large language models (LLMs), and vision models (Left). We observe that the initial layer intrinsic dimension is significantly lower in dec-LLMs and vision leading to a higher proportion of off-maifold AEs and consequently better robustness unlike to enc-LLMs. Motivated by this insight, we incorporate this in a new manifold-aware AT paradigm, SMAAT. SMAAT proceeds in two phases. First, it searches for the layer $l^{*}$ with the lowest ID. Second, SMAAT trains the network adversarially by generating samples from $l^{*}$ as opposed to $l_{1}$, as classically. This results in, $\mathcal{O}\left(P\times(n-l^{*})\right)$ gain in run-time complexity per training iteration (Right)
We evaluated the effectiveness of SMAAT on robustifying (1) classifiers, (2) safety filters, and (3) retrievers in Retrieval-Augmented Generation (RAG). SMAAT achieved state-of-the-art (SOTA) runtime and robustness results on all tasks, while maintaining clean accuracy (generalization) comparable to standard training (Fig.1). Specifically, in sentiment and content classification tasks, SMAAT improved the robustness of enc-LLMs on AGNEWS [76], IMDB [36], and YELP [74] datasets by 8.6%, 15.7%, and 28.8% for BERT [15] and by 6.0%, 5.8%, and 19.0% for RoBERTa [34], respectively. For safety filtering, SMAAT accurately identified harmful prompts generated by the GCG attack [81], achieving 97-100% accuracy. In RAG experiments, SMAAT significantly enhanced the robustness of the Contrevier model [23] on the RAG setup, achieving over 80% robustness against poisoning attacks [78]. Furthermore, SMAAT required only about 25-33% of the GPU time compared to the standard AT.
In summary, the major contributions of our work are:
  1. 1.
    SMAAT, a novel adversarial training algorithm that leverages the intrinsic dimensionality across layers of foundational models to control robustness and generalization.
  2. 2.
    An explanation for the discrepancy in the robustness and generalization trends in foundational models.
  3. 3.
    Comprehensive experiments demonstrating enhancements in robustness and scalability, while maintaining accuracy in classification and retrieval tasks compared to standard adversarial training.
2 Related Work
Adversarial Training (AT) aims at robustifying a model against AEs which are imperceptibly perturbed inputs that can lead to incorrect predictions. Formally, AT is framed as a robust optimization program. Specifically, it seeks the optimal parameters $\theta^{*}$ of a classifier $f_{\theta}(x)$ that are robust to perturbations $\delta$ within a norm ball, i.e.,
$\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\big{[}\max_{\|\delta\|\leq\epsilon}\ell(f_{\theta}(x+\delta),y)\big{]},$ (1)
where $\ell$ is a loss function (e.g., cross-entropy loss) and $\mathcal{D}=\{x,y\}_{i=1}^{|\mathcal{D}|}$ is the training data. It has been shown that the saddle-point problem can be solved reliably with SGD for the outer minimization and with project-gradient descent (PGD) for the inner maximization [38]. PGD starts at randomly initialized perturbations in the $\epsilon$-ball and iteratively applies a gradient ascent update followed by a projecting onto the $\epsilon-$ball $P$ times. $P$-step PGD AT is an order of magnitude more expensive than standard training for deep networks as the $P$-step PGD requires $P$ forward-backward passes through the network, while the standard SGD update requires only one.
Manifold-Based Defenses. The manifold conjecture stands as one of the most compelling explanations for the susceptibility of deep neural networks to adversarial samples [57, 21, 50]. The conjecture posits that data resides on a low-dimensional manifold within a high-dimensional representation space and that the deepnet model learns to approximate this manifold. Consequently, an off-manifold sample, deviating from this foundational manifold, leads to undefined behavior of the model. This conjecture has inspired a novel line of defenses against adversarial attacks on images [45, 39, 53, 47] and text [40]. These methods approximate the data manifold and, during testing, project samples onto this manifold to either detect or correctly classify. Differently, we leverage the manifold conjecture during training to improve both robustness and scalability.
Robustness and Generalization Trends in AT. AT is recognized for enhancing model robustness in both vision and enc-LLMs [73, 4]. While this improvement in robustness comes at the cost of increased generalization error in vision models [73], AT enhances generalization in enc-LLMs [4, 20]. There has been recently efforts relating robustness to OFM-AEs and generalization to ONM-AEs [56, 70]. We extend this line of research by connecting these trends to the intrinsic dimensionality of intermediate layers of a network.
Scalable AT. Different optimizations have been proposed to mitigate the cost of the PGD additional $P$ forward-backward passes for AEs generation, including (i) replacing multi-step PGD with a single-step FGSM [48]; (ii) omitting redundant computations during PGD [72]; (iii) combining FGSM AT with random initialization [68]. While these approaches alleviate the PGD overhead, they also come with limitations. As a solution, FreeLB [80] proposes accumulating model parameter gradients over multiple batches. In our work, we achieve scalability through an alternative method, i.e., by leveraging the manifold conjecture to perturb intermediate layers and hence reduce the length of backward-forward chains in PGD.
ID Estimation. ID corresponds to the minimum number of variables needed in a maximum representation of data. Various techniques have been developed to estimate ID, including (i) Two-Nearest Neighbors (twoNN), which examines distances to the two nearest neighbors [19], (ii) Singular Value Decomposition (SVD), used for projection error estimation by decomposing a matrix into singular values [54], (iii) Maximum Likelihood Estimation (MLE), which maximizes the likelihood of observed distances [43], and (iv) Correlation Dimension, which estimates ID based on the scaling behavior of pairwise distances [22]. In our work, we utilize twoNN for ID estimation due to its robustness to non-linearity of the data manifold. Specifically, it analyzes the ratio $\mu$ of distances from each point to its two nearest neighbors. The empirical distribution function of $\mu$ is modeled using a Pareto distribution model, expressed as $f(\mu|I)=I\mu^{-(I+1)}$, where $I$ is the ID of the manifold [19]. However, since twoNN is not applicable for estimating the OFM/ONM proportion of AEs, we utilize the SVD projection error estimation approach for this purpose.
3 Notation
Consider a deep neural network classifier $f_{\theta}$ with $n$ layers and parameters $\theta\!=\!\left\{\theta^{(i)}\right\}_{i=1}^{n}$. We denote by $f_{\theta^{[i,j]}}$ the nested transformation spanning layers $l_{i}$ to $l_{j}$, i.e., $f_{\theta^{[i,j]}}\!=\!f_{\theta^{(i)}}\!\circ\!\cdots\!\circ\!f_{\theta^{(j)}}$, and by $\bar{f}_{\theta^{[i,j]}}$ the transformation $f_{\theta^{[i,j]}}$ normalized to have zero mean and unit variance. Also, we denote by $\mathcal{\bm{X}}\in\mathbb{R}^{d\times|\mathcal{D}|}$ and $f_{\theta^{[1,l]}}(\mathcal{\bm{X}})\in\mathbb{R}^{d_{l}\times|\mathcal{D}|}$ the matrices obtained from stacking the input samples $\{x_{i}\}_{i=1}^{|\mathcal{D}|}$ and their transformations $\left\{f_{\theta^{[1,l]}}(x_{i})\right\}_{i=1}^{|\mathcal{D}|}$ vertically, for a given dataset $\mathcal{D}\!=\!\{x_{i},y_{i}\}_{i=1}^{|\mathcal{D}|}$.
The data manifold is described by the distribution of samples $\{x_{i}\}_{i=1}^{|\mathcal{D}|}$, i.e., samples in the high-density regions are on-manifold and samples in the low-density regions are off-manifold. Similarly, the $l^{\text{th}}$-layer manifold is characterized by the density of the normalized transformed samples, i.e., $\{\bar{f}_{\theta^{[1,l]}}(x_{i})\}_{i=1}^{|\mathcal{D}|}$. While these densities are unknown in practice, for simplicity we approximate the data and $l^{\text{th}}$-layer manifolds via the eigenspaces of the covariance matrices $\mathcal{\bm{X}}\mathcal{\bm{X}}^{T}$ and $\bar{f}_{\theta^{[1,l]}}(\mathcal{\bm{X}})\bar{f}_{\theta^{[1,l]}}(\mathcal{\bm{X}})^{T}$, respectively. We denote by $(U_{l},\Sigma_{l})$ the eigenspace of the $l^{\text{th}}$ layer, where $U_{l}\!\in\!\mathbb{R}^{d_{l}\times d_{l}}$ and $\Sigma_{l}\!\in\!\mathbb{R}^{d_{l}\times d_{l}}$ are the matrices of eigen-vectors and eigen-values, respectively. Since the eigenvectors $U_{l}$ are orthogonal (covariance matrices are symmetric), after normalization, the projection error on $k\in[1,d_{l}]$ eigenvectors is:
$e_{l}^{k}(x)=\bar{f}_{\theta^{[0,l]}}(x)-U_{l}^{k}{U_{l}^{k}}^{T}\bar{f}_{\theta^{[1,l]}}(x).$ (2)
The eigen-based data manifold of the $l^{\text{th}}$ layer is denoted by $k_{l}$ and is defined as the number of eigenvectors with the largest eigenvalues necessary for a reconstruction error $\gamma$ of the original inputs, i.e., $k_{l}\!=\!\min_{k}\left\{k\in[1,d_{l}]\Big{|}\sum_{x\in\mathcal{D}}\left\|e_{l}^{k}(x)\right\|_{2}\leq\gamma\right\}.$
An AE is $(\gamma,l)$-off-manifold, i.e., $(\gamma,l)$-OFM-AEs, if the projection error of its image at the input of the $l^{\text{th}}$ layer is larger than $\gamma$. Similarly, we define AEs as on-manifold (OMF-AEs) if $e_{l}^{k}\leq\gamma$.
The Intrinsic Dimension (ID) of the layer, estimated using the TwoNN method, is denoted as $I_{l}$. Furthermore, we denote the normalized ID of the layer by $I_{l}^{d}$ which is computed by dividing $I_{l}$ by $d_{l}$.
(a) BERT
(b) RoBERTa
(c) Vision Models
(d) dec-LLMs
Figure 3: The correlation between OFM/ONM AEs ratio (row 1) and the normalized ID (row 2). We use the average projection error ($e_{l}^{k}$) of AEs as a proxy for estimating the OFM/ONM AEs ratio and twoNN for ID estimation across layers of models including (a) BERT and (b) RoBERTa on AGNEWS, IMDB, and YELP datasets, (c) vision models on Cifar-10 dataset, and (d) dec-LLMs on HelpSteer dataset. Results show that enc-LLMs (BERT, RoBERTa) have decreasing ID trends unlike vision models and dec-LLMs that have increasing ID trends. These ID trends are reflected on reconstruction errors which show enc-LLMs have higher OFM/ONM AEs ratio at higher layers while vision models and dec-LLMs have at the initial layer.
4 Approach
4.1 Motivation and Overview
Our study is motivated by understanding the inherent difference in the generalization and robustness trends between adversarially trained vision and enc-LLMs. Specifically, (1) robustness improvement is consistently more significant in vision models [38, 4], and (2) generalization deteriorates in vision models, while it remains unchanged or even improves in enc-LLMs [61, 80].
We hypothesize that the underlying phenomenon is related to the distribution of OFM-AEs across the layers. Specifically, we speculate that in the first layer where conventional AT is applied, AEs tend to be more on-manifold for enc-LLMs, whereas they are more off-manifold for vision models and dec-LLMs. This would explain the results on generalization and robustness trends since it is established that higher proportions of on/off-manifold AEs lead, respectively, to better generalization/robustness [35, 55, 2, 31].
We study the relationship between the ID and the proportion of OFM/ONM AEs across different layers of various deepnet models, including two enc-LLMs (BERT and RoBERTa), three vision models (ViT [17], DeiT [58], VGG-13 [52]), and three dec-LLMs (LLaMA-2-7b [59], Vicuna [77], Mistral-7b [24]) on several dataset (refer to Appendix B for details). Our observation in the first row of Figure 3 shows that, for two enc-LLMs (BERT and RoBERTa), the proportion of OFM-AEs increases with the layer depth, in contrast to the trend observed in vision models and dec-LLMs. The empirical results confirm that normalized ID, $I_{l}^{d}$, follows the OFM/ONM ratio of AES since if the embedding space of the layer, $d_{l}$, is much greater than the ID of the layer, $I_{l}$, there is a higher probability for the samples to be OFM-AEs. In line with this, while $I_{l}^{d}$ decreases for the enc-LLMs, it increases for vision models [^2] Our findings are consistent with earlier studies that reveal a ”hunchback shape” —initially increasing and then decreasing– in the ID behavior of vision models [5]. As convolutional neural networks decrease in dimensionality $d_{l}$, the normalized ID, $I_{l}^{d}$, follows an increasing trend. and dec-LLMs. We also observe that estimated IDs and measured projection errors show a strong correlation across all models, with very few exceptions. (Refer to Appendix B for a further discussion on this issue.)
Inspired by these findings, we propose an approach for improving robustness over standard AT, via perturbing an intermediate layer $l^{*}$ with a low IDs. We show that our AT approach, Scalable Manifold Aware Adversarial Training (Fig. 2), significantly improves scalability by a factor $\mathcal{O}\Big{(}P(n-l^{*})\Big{(}\max\!\Big{(}\{d_{l_{i}}|l_{i}\in[1,l^{*}-1]\}\Big{)}-d_{l^{*}}\Big{)}\Big{)}$ for an $n$-layered deepnet under a $P$-step PGD attack.
4.2 SMAAT: Scalable Manifold Aware Adversarial Training
We propose SMAAT an adversarial training algorithm that aims at improving the robustness and scalability of AT by generating AEs in the layer leading to the highest proportion of OFM instead of from the input layer as classically. Specifically, given a pretrained model $f_{\theta}$, SMAAT intentionally generates a higher proportion of OFM-AEs to enhance robustness. Formally, SMAAT solves the augmented AT objective:
$\min_{\theta}$ $\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\max_{\|\delta\|\leq\epsilon}\ell(f_{\theta}(x+\delta),y)\right]$ (3)
[…]
While the manifold conjecture refers to input space AEs, we relax it to encompass intermediate layers as well, i.e., we chose the perturbation $\delta$ that results in an on-/off-manifold AE at any layer across the deepnet. Intuitively, AEs that are off or on the transformed data manifold in any layer also affect robustness. The relaxed objective becomes:
$\min_{\theta}$ $\mathbb{E}_{(x,y)\sim\mathcal{D}}\big{[}\max_{\|\delta\|\leq\epsilon}\ell(f_{\theta}(x+\delta),y)\big{]}$ (4)
[…]
Note that solving the above objective using the method of Lagrangian Multipliers is possible, but it would require an approximation of the manifold to characterize OFM-AEs. Such approximations are either computationally expensive (e.g., GAN [69, 45]) or overly simplistic (e.g., eigenbasis [70]). SMAAT uses an alternative approach to find OFM-AEs by perturbing the layer with the lowest ID without the need for manifold approximations. Particularly, SMAAT solves:
$\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big{[}\max_{\|\delta_{l^{*}}\|\leq\epsilon_{l^{*}}}\ell\Big{(}f_{\theta^{[l^{*},n]}}\Big{(}f_{\theta^{[1,l^{*}]}}(x)+\delta_{l^{*}}\Big{)},y\Big{)}\Big{]}.$ (5)
Here, $\delta_{l^{*}}$ is the perturbation at the $l^{*}$-th and $\epsilon_{l^{*}}$ is the perturbation budget. $l^{*}$ is the layer with more OFM-AEs composition (lowest ID as assumed in Sec. 4.1). We additionally choose $l^{*}$ to correspond to the layer with the highest index as this would lead to better scalability (shorter Backprop chains). i.e.,
$l^{*}=\max_{l}\Big{\{}l\in[1,n]\Big{|}I_{l}^{*}\leq I_{i}\forall i<l\Big{\}}.$ (6)
Complexity. Computing the layers IDs and searching the optimal layer $l^{*}$ to perturb are done once per model and per task and incur marginal overhead. For the AT part, when a $P$-step PGD attack is used, this results in $P$ forward-backward passes with length $(n-l^{*}+1)$ instead of $n$. The run-time of every forward/backward pass depends on the layer dimensionality, i.e., $O(d_{l-1}\times d_{l})$ for the $l^{\text{th}}$ layer. Overall, the complexity of one SMAAT forward-backward is $\mathcal{O}((n-l^{*}+1)\max_{l\in[l^{*},n]}(d_{l})^{2}))$. As a result, SMAAT is more efficient than standard AT by a factor of $\mathcal{O}(P\times l^{*}\cdot(\max_{i\in[1,n]}{(d_{i})^{2}}-\max_{j\in[l^{*},n]}{(d_{j})^{2}}))$. In the case of BERT and ROBERT, $l^{*}$ is equal to $n$. The total run-time can be simplified to $\mathcal{O}(P\cdot(d_{n})^{2})$ where $d_{n}$ is the number of classes. Typical enc-LLMs tasks consist of less than five classes [62] which makes the factor of classes small enough to be negligible. Hence, SMAAT enhances the efficiency of the AEs generation process by a factor of $l\cdot O(\max(d_{l}))$. This improvement practically eliminates the cost of the AE generation process.
5 Experiments
We assess SMAAT for its ability to enhance the robustness of various components associated with the use of a language model, aiming to highlight its practical impact. These include: (i) the model itself, particularly in relation to classification tasks; (ii) the retriever model within the Retrieval-Augmented Generation (RAG) setting; (iii) the safeguarding models employed for moderating content produced by generative models. Our robustness evaluation considers four different attacks: (i) word substitution attacks [44, 25, 30], (ii) adaptive attacks [60], (iii) Greedy Coordinate Gradient (GCG) attack [81], and (iv) corpus poisoning attack [78]. We also include a run-time analysis to demonstrate scalability of our method.
5.1 Robustifying Classification Models
Base models. We robustify two base models, BERT-base-cased and RoBERTa-base-cased. Tasks. Considered tasks sentiment clasification (IMDB [36] and YELP [74]) and topic classification (AGNEWS [76]). Attacks. We assess the robustness of models against three input space attacks including PWWS [44] (synonym based), TextFooler [25] (neighbor based), and BERT-Attack [30] (masked language model based). All attacks are conducted using the TextAttack framework [41] and following the settings introduced by Li et al. [32].
Baselines. We compare SMAAT to several baselines including standard (non-adversarial) training, and six baselines from three families of defenses: (1) Input space AT (ASCC [16], DNE [79]), (2) Embedding space AT (FreeLB++ [32]), and (3) Certified defenses (SAFER [71], TMD [40], RSMI [40]). FreeLB++ focuses on scalability and robustness by minimizing the number of PGD steps and applying AT in the initial layer. TMD leverages manifold features by projecting samples back to the manifold in the last layer.
Implementation details. In both base models, $l^{*}=l_{n}$, i.e., we generate AEs in the last layer. Specifically, we perturb the [CLS] embeddings before the classifier layer. We freeze all layers before $l^{*}$ and adversarially train the layers following it for ten epochs with varying attack strength $\epsilon\in[0.1,0.8]$ with learning rate of $0.1$. Further details are presented in Appendix C.
Table 1: Robustifying sentiment classifiers on AGNEWS, IMDB, and YELP datasets. We report the clean accuracy (CA) and the robust accuracy under attack (AUA) with PWWS (PW), TextFooler (TF), and Bert-Attack (BA), along with the average robust accuracy (AR) across these three attacks. The best performance is bolded.
Model Defense AGNEWS IMDB YELP
CA AUA CA AUA CA AUA
PW TF BA AR PW TF BA AR PW TF BA AR
BERT Standard 94.5 36.9 28.1 37.5 34.2 92.2 15.0 5.8 5.4 8.7 97.0 12.2 6.5 5.3 8.0
ASCC 91.6 32.8 31.4 32.1 32.1 88.5 15.1 12.4 11.2 12.9 91.5 19.4 15.7 12.2 15.8
DNE 94.1 34.0 33.6 52.3 40.0 90.0 25.7 23.0 20.6 23.1 94.0 33.3 31.2 43.8 36.1
FreeLB++ 95.1 47.9 51.5 41.8 47.1 93.2 12.5 45.3 39.9 32.6 95.6 19.3 8.8 3.7 10.6
SAFER 94.4 39.3 35.5 42.3 39.0 92.3 41.4 39.1 30.7 37.1 95.4 29.8 25.8 23.7 26.4
TMD 94.3 70.0 50.0 55.2 58.4 92.2 38.7 44.2 33.7 38.9 95.2 36.8 40.9 28.6 35.4
RSMI 92.7 76.1 63.2 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA 92.2 58.7 56.4 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA 95.4 45.3 52.3 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA
SMAAT (Ours) 94.6 73.5 72.2 74.7 73.5 92.2 63.6 77.9 60.8 67.4 97.0 77.1 77.9 72.8 75.9
RoBERTa Standard 94.7 30.6 23.9 37.1 30.5 94.0 8.7 2.1 0.6 3.8 97.9 23.1 14.9 9.0 15.7
ASCC 92.6 48.1 41.0 49.1 46.1 92.6 23.1 13.5 11.8 16.1 95.4 15.0 8.6 4.5 9.4
DNE 94.9 58.0 46.5 54.5 53.0 94.2 48.8 26.9 16.0 30.6 96.8 64.4 64.0 45.2 57.9
FreeLB++ 95.6 61.0 49.8 56.6 55.8 94.3 33.6 14.6 6.1 18.1 97.0 38.6 46.0 35.2 39.9
SAFER 94.6 68.9 49.3 46.1 54.8 93.9 52.8 47.1 40.6 46.8 96.6 65.6 67.9 48.3 60.6
TMD 95.0 68.3 54.0 56.7 59.7 93.3 60.5 66.8 51.6 59.6 96.6 68.9 70.9 51.0 63.6
RSMI 94.3 81.9 74.1 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA 93.0 76.2 73.4 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA 96.3 68.9 65.9 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA
SMAAT (Ours) 94.6 75.6 75.1 79.9 76.9 93.5 77.1 78.5 63.2 72.9 98.0 85.4 86.4 76.0 82.6
Robustness and generalization.
In Table 1, we report results on robustness and generalization. Note that generalization is high if the adversarially trained models result in higher clean accuracy (CA) than standard training (1st row). On average, SMAAT demonstrates superior robustness over all datasets: it achieves an improvement of 8.6%, 15.7%, and 28.8% over the best score for the BERT model, and 6.0%, 5.8%, and 19.0% for the RoBERTa model on the AGNEWS, IMDB, and YELP datasets, respectively. Note that FreeLB++, which perturbs the first layer, showed the best generalization in four out of six cases, as it produces mostly on-manifold examples (the first layer has a high dimension, as illustrated in Fig. 3) which is inline with the manifold conjecture. SMAAT maintains generalization in five out of six cases and only shows $0.5$ drop in performance in the case of RoBERTa with the IMDB dataset. Besides, SMAAT consistently outperforms TMD, even though it is a manifold based method differently from SMAAT, TMD estimates the manifold and projects the input samples onto it prior to classification. This might lead to inaccuracies in the manifold estimation. SMAAT on the other hand is more robust to such inaccuracies as it only leverage the ID for selecting the layer to perturb. From an attack perspective, RSMI demonstrates superior robustness against PWWS, a synonym-based word substitution attack, in both models. This phenomenon can be attributed to the behavior of the masked inference (MI) component in RSMI, which mimics a synonym-substitution based defense approach without requiring an explicit synonym set, as noted in the original paper [40].
Table 2: Run-time results on IMDB dataset. Mean and standard deviation are computed over ten runs.
Standard ASCC DNE FreeLB++ SAFER RSMI SMAAT
Training (min/epoch)
5.1 ±0.1 25.7 ±0.3 15.2 ±0.9 15.6 ±0.5 8.2 ±0.6 15.4 ±0.3 5.2 ±0.2
Inference (msec/sample)
2.4 ±0.1 41.4 ±0.2 4.0 ±0.2 2.4 ±0.0 2.4 ±0.0 5.6 ±0.4 2.4 ±0.1
Runtime efficiency. In Table 2, we provide details on the training time per epoch and inference time per instance for the IMDB dataset using BERT (RoBERTa has the same architecture). SMAAT has comparable efficiency to standard training and is on average, 3 times faster than standard AT during training. This is attributed to shorter backpropagation chains as AT in SMAAT is performed in the last layer ($l^{*}=l_{n}$). Note that FreeLB++, which inject noise in the first layer, remains inefficient even when PGD is replaced with FGSM [80, 32], as a full-depth backpropagation chain is still required. Certified defense baselines (SAFER, RSMI) and input space attacks (ASCC, DNE) are also time-consuming as they either require mapping samples into the manifold or performing an extensive search over word substitutes.
Figure 4: Effect of applying AT on generalization/robustness when applied to middle layers of BERT on the YELP dataset against TextFooler attack. Model robustness improves as adversarial training is applied to higher layers while generalization remains similar.
AT in intermediate layers. We experimentally show that, under decreasing ID behavior, AT in higher-layers yields increased robustness with increasing layer index. Thus, we apply AT to layers $l_{i}$ with $i\in\{0,2,4,6,8,10,11,\text{and }12\}$. We keep the number of PGD steps to 3 and the number of training epochs to 5 across layers. We report results on AT of BERT on the YELP dataset in Fig. 4. As expected, robustness (under TextFooler attack) increases with the layer index $i$ without affecting generalization. This observation aligns with our findings in Figure 3. The rise in OFM-AE ratio corresponds to an increase in robustness as we progress to higher layers.
Additional robustness evaluation of SMAAT on language understanding benchmarks, including GLUE and advGLUE are reported in Appendix D. Also, to emphasize that in the general case AT at the last layer is not sufficient for improving robustness, we conduct experiments as detailed in Appendix E.
5.2 Robustifying Safety Filters in Dec-LLMs
Table 3: Robustifying safety filters based on dec-LLMs on AdvBench and HH_RLHF datasets. We report clean accuracy (CA), and robust accuracy under attack (AUA) which indicates the percentage of harmful prompts augmented with adversarial suffices accurately classified as harmful.
Dataset Model BERT RoBERTa DistilBERT
CA AUA CA AUA CA AUA
AdvBench Standard 100 0 100 0 99.6 0
FreeLB++ 99.6 0 100 0 99.6 0
SMAAT 100 100 100 99.2 99.2 97.5
HH-RLHF Standard 95.4 26.4 94.8 0.3 98.7 6.5
FreeLB++ 98.6 48.5 98.6 34.0 98.5 34.1
SMAAT 95.4 51.6 94.8 96.8 98.5 39.5
Table 4: Robustifying retriever models of RAG on the Natural Questions dataset. We report Recall@10 (R@10) and Recall@100 (R@100) on the clean corpus for generalization. Robust recall (RR) is measured by how many samples are selected without adversarial passages in the top-k passages, with R@10 and R@100 corresponding to the top-10 and top-100 passages, respectively. During attacks, 10 and 50 adversarial passages are created, denoted as (N=10) and (N=50), respectively.
Model Generalization RR (N=10) RR (N=50)
R@10 R@100 R@10 R@100 R@10 R@100
Standard 36.4 79.7 46.0 22.2 26.1 7.2
FreeLB++ 39.6 81.8 75.2 60.7 51.7 32.2
SMAAT 34.9 79.5 99.5 97.1 85.7 73.5
The broad attack surface of LLMs has compelled model owners to implement solutions that extend beyond the safety alignment of a model. Notably, this includes content moderation filters that verify the harmlessness of a model’s inputs and outputs [27, 9]. These filters are essentially text classifiers, often based on lightweight enc-LLMs to minimize overhead and preserve the high responsiveness of dec-LLMs [27]. The robustness of safety filters is also a concern, as they too can be vulnerable to attacks. Adversarial attacks. The Greedy Coordinate Gradient (GCG) attack [81], which generates adversarial suffixes that are appended to the input prompts of a language model to bypass safety alignment, proves effective across both encoder-based and decoder-based architectures. Base models. We evaluate SMAAT’s effectiveness in improving the robustness of BERT, RoBERTa, and DistilBERT models when employed as safety filters. Dataset. Tests are performed on two datasets: the AdvBench dataset [81] and the Helpfulness-Harmfulness Human Preference dataset (HH-RLHF) [7]. We use GCG attack. Metrics. In our tests, we benchmarked the filter’s robustness by considering its generalization accuracy (ACC) for safe and harmful prompts, and robust accuracy under attack (AUA) by measuring the detection rate of harmful prompts augmented with adversarial suffixes. Further details are presented in Appendix F.
The results in Table 4 show that SMAAT significantly enhances the robustness of safety filters against the GCG attack across all models, while maintaining generalization accuracy compared to standard training. In line with our observation, FreeLB++ improves generalization as it applies AT in the first layer.
5.3 Robustifying Retriever Models of RAG
Baselines. We evaluate the robustness of retriever models within the Retrieval-Augmented Generation (RAG) framework under standard, FreeLB++ and SMAAT. RAG combines a retriever model, which identifies relevant passages from a large corpus, with a generator model that constructs answers based on the retrieved information. Adversarial attacks. Poisoning attacks [78], which manipulate the retrieval corpus by generating adversarial passages, are particularly critical in this setup. Base model. We use the Contriever model [23], fine-tuned on the Natural Questions (NQ) [29] dataset, as our retriever. Metrics. We use robust recall (RR) measured by how many samples are selected without adversarial passages in the top-k passages, with R@10 and R@100 corresponding to the top-10 and top-100 passages, respectively.
Results in Table 4 show that the standard model offers limited robustness against the poisoning attack, with a significant number of adversarial passages being retrieved. FreeLB++ showed some improvement, reducing the number of adversarial passages retrieved. However, SMAAT significantly enhanced robustness, demonstrating a dramatic reduction in the retrieval of adversarial passages compared to the other methods. In terms of generalization, FreeLB++ yields the best results since it applies to AT in the initial layer (more ONM-AEs) while SMAAT and standard training have similar performance. In terms of generalization, FreeLB++ yields the best results as it applies AT in the first layer (resulting in more ONM-AEs), while both SMAAT and standard training exhibit similar performance.
6 Conclusion
In this paper, we introduce a manifold-aware approach designed to enhance the scalability, and robustness of adversarially trained deep neural networks while maintaining generalization. Our approach involves generating adversarial examples in higher layers with the lowest intrinsic dimensionality (ID). This strategy enables faster synthesis with shorter backpropagation chains and enhances robustness by capturing a higher proportion of off-manifold AEs. We achieve state-of-the-art performance in both runtime efficiency and robustness, while maintaining comparable levels of generalization to standard training. There are several avenues for future work including (i) investigate approaches for joint optimization of both generalization and robustness through perturbations in the intermediate layers, i.e., via controlling the on-off manifold AEs ratio (ii) study the impact of off-manifold AEs on out-of-distribution detection.
Appendix A Supplementary Material
In this paper, we propose SMAAT, an adversarial training algorithm that not only optimizes for better robustness and generalization but also enhances scalability. Specifically, SMAAT leverages the manifold conjecture, which posits that off-manifold adversarial examples (OFM-AEs) lead to better robustness while on-manifold adversarial examples (ONM-AEs) enhance generalization. To achieve this, SMAAT perturbs the layer with the lowest intrinsic dimension (ID). Intuitively, such a layer is expected to yield the highest proportion of off-AEs. Formally, SMAAT solves the following program:
$\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big{[}\max_{\|\delta_{l^{*}}\|\leq\epsilon_{l^{*}}}\ell\Big{(}f_{\theta^{[l^{*},n]}}\Big{(}f_{\theta^{[1,l^{*}]}}(x)+\delta_{l^{*}}\Big{)},y\Big{)}\Big{]}.$ (7)
We show that $l^{*}$ corresponds to the last layer in encoder-based LLMs (enc-LLMs) and the first layer in decoder-based LLMs (dec-LLMs). This explains the difference in robustness and generalization trends between vision models and enc-LLMs [55, 48, 80]. Specifically, in vision models, improvements in robustness are often accompanied by a decrease in generalization. In contrast, for enc-LLMs, improvements in generalization are achieved with lesser gains in robustness.
SMAAT has the advantage of being significantly more runtime-efficient than classical adversarial training, by a factor $\mathcal{O}\Big{(}P(n-l^{*})\Big{(}\max\!\Big{(}\{d_{l_{i}}|l_{i}\in[1,l^{*}-1]\}\Big{)}-d_{l^{*}}\Big{)}\Big{)}$ for an $n$-layered deepnet under a $P$-step PGD attack where $d_{l}$ is the dimensionality of the layer $l$. Empirical results demonstrate that SMAAT leads to better robustness while maintaining comparable accuracy to standard training. Applications considered include robustifying (1) sentiment classifiers, (2) safety filters in decoder-based models, and (3) retriever models in the Retrieval Augmented Generation (RAG) setup.
The supplementary material is organized as follows:
  • Appendix B: Intrinsic Dimension Estimation
  • Appendix C: Additional Results of Robustifying Sentiment Classifiers
  • Appendix D: Additional Results of Language Understanding Benchmarks
  • Appendix E: Additional Results AT at the Last Layer
  • Appendix F: Additional Results of Robustifying Safety Filters In Decoder Based LLMs
Appendix B Intrinsic Dimension Estimation
In the following, we provide additional details on the experimental setup on off-manifold ratio calculation and ID estimation in Figure 3. Experimental setup. We conduct experiments on two enc-LLMs (BERT and RoBERTa), across three different datasets (AGNEWS, IMDB, YELP). Additionally, we evaluated three vision models (ViT [17], DeiT [58], VGG-13 [52]) on the CIFAR-10 [26] dataset and three LLMs (LLaMA-2-7b [59], Vicuna [77], Mistral-7b [24]) on the HelpSteer dataset  [66]. We utilize the train split of the datasets to estimate the ID and eigenspace of the layers. AEs are generated using TextFooler for LMs and PGD for vision models on the respective test splits of the datasets. For LLMs, AEs are generated using the Greedy Coordinate Gradient (GCG) attack [81] on the harmful prompt datasets released in the same paper. In our calculation, we use CLS embedding for BERT, RoBERTa, ViT, and DeiT models; last token embedding for LLMs; and conventional layer output for VGG-13.
Discussion of results. We present the average projection error, $e_{l}^{k}$, of each layer on the first row of Fig. 3. Results in Fig. 3(a,b) indicate that for LMs, the average $e_{l}^{k}$ monotonically increases, suggesting that examples become more off-manifold at the latest layers, consistent with our hypothesis. Conversely, for vision models and dec-LLMs in Fig. 3(c,d), we observe the opposite characteristic with a lower average $e_{l}^{k}$ at the latest layers as expected. The only unexpected behavior observed in vision models and dec-LLMs, except for VGG-13 and Mistral, is that they exhibit the highest off-manifold ratio ($e_{l}^{k}$) in the initial layers rather than the first layers. For transformer models, the CLS/last token embedding starts with a specific value and gradually evolves to represent the sentence. This means it takes time for the effect of AEs to manifest in the CLS/last token embedding. In contrast, since the VGG model’s representation is directly calculated from the input, the highest off-manifold ratio is obtained at the first layer.
We hypothesize that this phenomenon can be explained by the ID of the layers. Specifically, if $I_{l}\ll d_{l}$, AEs tend to be off-manifold. To validate this hypothesis for the same models and dataset, we measure $I_{l}$ using the twoNN method and normalize it with $d_{l}$. The normalized $I_{l}$ can be seen in the second row of Fig. 3. In line with our hypothesis, while it decreases for the LMs (Fig. 3(a,b)), it increases for vision models and dec-LLMs(Fig. 3(c,d)).
Appendix C Additional Results of Robustifying Sentiment Classifiers
In the following, we provide more details about the experiments on robustifying sentiment classifiers.
Datasets. We evaluate SMAATon three datasets: AG-News Corpus (AGNEWS) [75], Internet Movie Database (IMDB) [37], and Yelp Review Polarity (YELP) [74]. The AGNEWS dataset contains over 120000 samples, each belonging to one of the four labels: World, Sports, Business, Sci/Tech. The IMDB dataset contains 50000 data samples of movie reviews with binary labels for negative and positive sentiments. The YELP dataset contains nearly 600000 samples of highly polar Yelp reviews with binary labels. However, due to limitations in computing resources, we only use a subset of 63000 samples of the YELP dataset. In addition, we randomly sample 10% of the training set for validation in all datasets. For testing, we use a subset of 1000 test samples from each dataset, following previous work practices. The AGNEWS dataset contains over 120k samples, categorized into four classes: World, Sports, Business, and Sci/Tech. The IMDB dataset consists of 50k movie reviews, each labeled with binary sentiments (positive or negative).
Base model. We employed the $\text{BERT}_{\text{base-cased}}$ [15] and $\text{RoBERTa}_{\text{base-cased}}$ [34] models in our experiments. To conduct the evaluations, we utilize the fine-tuned models provided by TextAttack from HuggingFace for all datasets, except for the RoBERTa base model fine-tuned on YELP dataset. For the YELP dataset, we created a fine-tuned RoBERTa model for 2 epochs with a learning rate of $051e-$ 05 and a batch size of 32.
Adversarial Attacks. which include the following constraints: (1) The maximum percentage of modified words is set to $0.3$ for AGNEWS, $0.1$ for IMDB and YELP datasets, respectively. (2) For word replacement, a maximum of $50$ candidates are considered for each word. (3) The semantic similarity, measured using the Universal Sentence Encoder [11], between the original input and the generated adversarial example must exceed $0.84$. PWWS uses word synonyms, TextFooler applies nearest neighbor search in counter-fitting embeddings [42], and BERT-Attack utilizes BERT masked language model to generate candidate words.
Baselines. For input space adversarial training, we consider Adversarial Sparse Convex Combination (ASCC) [16] and Dirichlet Neighborhood Ensemble (DNE) [79]. These methods model the perturbation space as the convex hull of word synonyms. ASCC incorporates an entropy-based sparsity regularizer to capture word substitution geometry more effectively, while DNE expands the convex hull to encompass all synonyms of each word’s synonyms, combining Dirichlet sampling and adversarial training to enhance model robustness. In our investigation of embedding space adversarial training which recognized as the most impactful technique for enhancing generalization [33], we conduct a thorough analysis of FreeLB++ [32] (employs gradient-guided perturbations centered around the most susceptible data points). For certified defenses, we evaluate SAFER [71], TMD [40], and RSMI [40]. SAFER constructs a set of randomized inputs by performing random synonym substitutions and using the statistical properties of predicted labels to certify robustness. TMD employs infoGAN [12] to project adversarial examples to the data manifold in the last layer to address the manifold issue. RMSI combines these ideas by applying importance-based masking to tokens and leveraging randomized smoothing in each layer.
Implementation details. To train the last layer of $f_{\theta}$ with adversarial samples, we create adversarial samples using 5-step PGD attacks. During training, we use epsilon values of $0.1$, $0.1$, and $0.8$ for the YELP, AGNEWS, and IMDB datasets, respectively, for the BERT models. For the RoBERTa models, we employ epsilon values of $0.1$, $0.6$, and $0.03$. All models are trained $10$ epochs with a learning rate of $0.1$.
Appendix D Additional Results of Language Understanding Benchmarks
To comprehensively evaluate SMAAT’s performance against a broader spectrum of textual adversarial attacks, we employ the GLUE and AdvGLUE benchmarks. The GLUE benchmark [63] is a comprehensive evaluation suite featuring seven diverse NLP tasks to assess model performance. The AdvGLUE benchmark [64] is an extension of GLUE, incorporating 17 distinct textual adversarial attacks, covering word-level transformations, sentence-level manipulations, and human-written AEs. This extension ensures a thorough evaluation encompassing various adversarial linguistic phenomena. For our assessment, we employ the evaluation sets of four datasets across three different tasks: Sentiment Analysis (SST-2), Duplicate Question Detection (QQP), and Natural Language Inference (QNLI, RTE).
In our evaluation, we compare SMAAT against standard BERT and RoBERTa models [^4] We use the fine-tuned models available from https://huggingface.co/JeremiahZ, as well as their FreeLB++ incorporated versions. In the case of SMAAT, we conducted a grid search for the learning rate, ranging from $0.1$ to $0.001$, and the $\epsilon$ value, ranging from $0.8$ to $0.01$, using 3-PGD steps As shown in Table D, SMAAT demonstrates a robustness improvement of 5.6% and 2.6% for BERT, and 12.1% and 2.5% for RoBERTa, compared to the standard and FreeLB++ models, respectively, while maintaining similar generalization performance.
Table 5: Average accuracy from the standard training, FreeLB++ and SMAAT on GLUE and AdvGLUE datasets. Results clearly demonstrate that SMAAT enhances model generalization (GLUE results) and robustness (AdvGLUE results).
{NiceTabular}
c|ccc|ccc Dataset BERT   RoBERTa
Standard FreeLB++ SMAAT (Ours) Standard FreeLB++ SMAAT (Ours)
GLUE 85.9 86.3 86.3 89.3 89.6 89.7
AdvGLUE 39.5 42.5 45.1 27.5 37.1 39.6
Table 6: Hyperparameters in experiments on robustifying the safety filters on decoding LLMs experiments. Learning rate (lr) and $\epsilon$ values for SMAAT and FreeLB++ methods.
Dataset FreeLB++ SMAAT
BERT RoBERTa DistilBERT BERT RoBERTa DistilBERT
lr $\epsilon$ $lr$ $\epsilon$ $lr$ $\epsilon$ lr $\epsilon$ $lr$ $\epsilon$ $lr$ $\epsilon$
AdvBench 5e-3 0.05 5e-3 0.05 1e-2 0.01 1e-4 1e-3 1e-4 1e-3 1e-4 1e-4
HH-RLHF 5e-3 0.5 53-3 0.05 5e-4 0.05 1e-4 5e-4 1e-4 1e-3 1e-4 5e-4
Appendix E Additional Results AT at the Last Layer
For vision models and decoder-based LLMs (dec-LLMs), the observed intrinsic dimension (ID) characteristics suggest that $l^{*}$ corresponds to the initial layer. This finding is in line with YOPO [72], which emphasizes the critical role of the initial layers for adversarial training (AT) in vision models. For the sake of completeness, we provide additional experiments on applying AT exclusively to the last layer of the vision and dec-LLMs models. For vision models, we train a Vision Transformer (ViT) model on the CIFAR-10 dataset. For dec-LLMs, we train the LLAMA-2-7b model with 8-bit quantized LoRA [14] on the SST-2 dataset using the SFTTrainer library from HuggingFace [1].
Figure 5: Robust accuracy obtained from standard training, AT in the input layer ($l_{1}$) and AT at the last layer. Results demonstrate that applying AT at the last layer doesn’t consistently lead to improved robustness. This is inline with our finding.
Our results from different types of models in Fig. 5 show that while applying adversarial training (AT) at the last layer improves robustness for enc-LLMs, it performs similarly to standard training for vision and dec-LLMs. These results motivate the ID analyses of SMAAT.
Appendix F Additional Results of Robustifying Safety Filters In Decoder Based LLMs
The GCG attack is deliberately designed to bypass the safety alignment of LLMs by generating a response to a potentially harmful prompt through appending an adversarial suffix to the user’s input. This strategy leverages previous methodologies by using (1) an affirmative response tactic [67, 10] to direct the model’s output towards the attacker’s intended outcome (used for loss calculation) and (2) a mix of greedy and gradient-based discrete optimization [51] to pinpoint the most susceptible tokens. A key attribute of the GCG attack is its transferability, demonstrating that adversarial suffix designed for a specific prompt on one model can successfully affect a broad range of other models.
A suggested defense against the GCG attack involves incorporating a lightweight binary classifier model designed to identify harmful prompts [27]. It is important to note, however, that these classifiers can still be vulnerable to such attacks. To assess the effectiveness of SMAAT against the GCG attack, we conduct the attack by setting the suffix length to 20 tokens. The attack is generated over 50 iterations, with 100 trials per iteration. Rather than leveraging the attack’s transferability, we tailor an adversarial suffix for every individual prompt and model to enhance the attack’s impact. Our assessment is conducted using two datasets, AdvBench [81] and HH-RLHF [7], and involves three models: BERT, RoBERTa, and DistilBERT. AdvBench features 640 training prompts (320 harmful, 320 safe) and 240 test prompts (120 harmful, 120 safe), whereas HH-RLHF includes approximately 44K harmful and 44K helpful training prompts, with 2.3K test samples for each prompt category. In all cases, standard models are trained over 5 epochs with a learning rate of $1e^{-5}$. Table 6 details the training hyperparameters for SMAAT.